One morning, Justin drove 2 hours before stopping to eat lunch at Chick-fil-A. After lunch, he increased his speed by 10 mph. If he completed a 260-mile trip in 7 hours of driving time, how fast did he drive in the morning?
Answer by ikleyn(52777)

.
Let x be his speed at the morning, in miles per hour.


Then his speed after lunch was (x+10) mph.


The equation for the total distance is


2x + 5*(x+10) = 260.


2x + 5x + 50 = 260


7x = 260-50 = 210


x = 210/7 = 30.


Answer.  His speed in the morning was 30 mph.

Solved.

-----------------

See introductory lessons on Travel and Distance problems
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.